### v1

Abstract
There is widespread concern about the negative impacts of social media feed ranking algorithms on political polarization. Leveraging advancements in large language models (LLMs), we develop an approach to re-rank feeds in real-time to test the effects of content that is likely to polarize: expressions of antidemocratic attitudes and partisan animosity (AAPA). In a preregistered 10-day field experiment on X/Twitter with 1,256 consented participants, we increase or decrease participants’ exposure to AAPA in their algorithmically curated feeds. We observe more positive outparty feelings when AAPA exposure is decreased and more negative outparty feelings when AAPA exposure is increased. Exposure to AAPA content also results in an immediate increase in negative emotions, such as sadness and anger. The interventions do not significantly impact traditional engagement metrics such as re-post and favorite rates. These findings highlight a potential pathway for developing feed algorithms that mitigate affective polarization by addressing content that undermines the shared values required for a healthy democracy.
keywords:
social media, polarization, feed ranking, field experiment
1 Main
Social media algorithms profoundly impact our lives: they curate what we see [1] in ways that can shape our opinions [2, 3, 4], our moods [5, 6, 7], and our actions [8, 9, 10, 11, 12]. Due to the power that these ranking algorithms have to direct our attention, many have expressed concern that they may be exacerbating negative societal outcomes [13, 14, 15, 16, 17]. Of particular concern is whether feed algorithms cause affective polarization—hostility toward opposing political parties [18, 19, 20]. If social media algorithms are causing affective polarization, they might not only bear responsibility for rising political incivility online [21] but also pose a risk to trust in democratic institutions [22]. In this case, isolating the algorithmic design choices that cause polarization could offer alternative algorithmic approaches.
Research investigating this question has so far yielded mixed results due to limited collaboration between platforms and researchers, frequent changes in feed-ranking algorithms, and the lack of strong incentives for platforms to address the issue [23]. A major hypothesized mechanism for how feed algorithms cause polarization is a self-reinforcing engagement loop: users engage with content aligning with their political views, the feed algorithm interprets this engagement as a positive signal, and the algorithm exposes even more politically aligned content to users, leading to a polarizing cycle. Some studies support this hypothesis, finding that online interactions exacerbate polarization [24], potentially because of the increased visibility of hostile political discussions [25], divisive language [26, 27, 28, 29, 30], and content that reinforces existing beliefs [31]. However, large-scale field experiments aimed at reducing polarization by intervening on the feed algorithm, for example by increasing exposure to out-party content, have found both a decrease [32] and increase [33] in polarization. Similarly, recent large-scale experiments on Facebook and Instagram found no evidence that exposure to in-party sources or a simpler reverse-chronological algorithm affected polarization and political attitudes [34, 35] during the 2020 U.S. election [36]. These mixed results have led to difficulty in identifying what, if any, algorithmic intervention might help.
We can distill the goals of these prior interventions to an alternative hypothesis: that feed algorithms cause affective polarization by exposing us to content that polarizes. An algorithm that upranks genuine political dialogue is less likely to polarize than an algorithm that upranks demagoguery. This hypothesis has been difficult to operationalize into interventions, making studies that intervene on cross-partisan exposure and reverse-chronological ranking attractive but more diffuse in their impact and thus more likely to observe mixed results. However, recent advances in AI large language models (LLMs) now provide an opportunity to create a ranking intervention that more directly targets the focal hypothesis [37]. We draw in particular on a recent large-scale field experiment that articulated eight categories of antidemocratic attitudes and partisan animosity as bipartisan threats to the healthy functioning of democracy [38]. We operationalize these eight categories into an AI classifier that labels expressions of these constructs in social media posts, does so as accurately as trained annotators agree with each other, and produces the desired depolarization effects in a lab setting on a fixed feed [39]. This real-time classification enables us to perform a scalable, content-based, real-time reranking experiment on participants’ own feeds in the field [40].
We conduct a preregistered field experiment on X (Twitter), the most used platform for political discourse in the US [41], that re-ranks participants’ social media content by either increasing or decreasing exposure to content that expresses these eight factors of antidemocratic attitudes and partisan animosity (AAPA). Consented participants (Sec. 6) installed a web extension and experienced this intervention through a real-time reranking of their feeds when they visited the X website over the course of a week. We measured the intervention’s effect on affective polarization [42] and emotional experience [43]. Compared to control conditions that did not rerank the feed, increased AAPA exposure in the ranking algorithm led to colder feelings toward the political outgroup, whereas decreased AAPA exposure led to warmer feelings. These changes also had an effect on participants’ levels of negative emotions—anger and sadness—as measured through in-feed surveys. Our intervention does not lead to statistically significant changes to traditional engagement metrics. However, because we observe higher engagement on political social media content than other content, the engagement-prioritizing nature of current feed algorithms is likely already amplifying AAPA content.
This work demonstrates not just causal evidence of feed algorithm polarization but a potential intervention to mitigate it by downranking content that violates shared democratic values.
2 Exposure to Antidemocratic Attitudes and Partisan Animosity
2.1 Antidemocratic Attitudes and Partisan Animosity (AAPA)
There are many viable theories as to which content, if amplified, could impact affective polarization. In this experiment, we draw specifically on prior work that articulates eight factors as problematic outcomes for healthy democratic functioning, referred to as “antidemocratic attitudes and partisan animosity” [38]: (1) partisan animosity, (2) support for undemocratic practices, (3) support for partisan violence, (4) support for undemocratic candidates, (5) opposition to bipartisanship, (6) social distrust, (7) social distance, and (8) biased evaluation of politicized facts (Sec. A.1). We refer to these eight factors collectively as AAPA (for Antidemocratic Attitudes and Partisan Animosity), and we hypothesize that increased or decreased exposure to AAPA content will cause corresponding changes to affective polarization. Our feed ranking intervention first identifies only political content in the feed, using a broad definition from the Pew Research Center [44] that includes mentions of officials and activists, social issues, news, and current events (Sec. B.3). We then score each political post from zero to eight based on the number of the eight AAPA factors expressed in its text, and we classify a post as AAPA if it reflects at least four of these eight factors. This scoring is performed by a large language model (Sec. B.4), which shows a high correlation with human judgment on identifying these eight factors in social media posts [39].
2.2 Experimental design
Figure 1: Overview of the field experiment on X. (A) Participants complete a pre-experiment survey to rate their feelings toward the opposing party and report their emotions. Then, they are randomly assigned to one of the two parallel experiments, increased or reduced AAPA content exposure, and further randomized into treatment and control groups. After 10 days, they complete a post-experiment survey. (B) Timeline of the experiment, which spans ten days. During the first three days, no intervention is applied to measure the participants’ baseline responses. After day three, participants in the treatment groups received increased or decreased AAPA exposure. In-feed surveys are occasionally shown throughout the baseline and the intervention period. (C) The interventions downranked or upranked AAPA posts in the participants’ feeds, depending on their treatment assignment. In-feed surveys are occasionally added to the feed to capture participants’ in-feed responses.
We designed two parallel experiments to measure the impact of reducing and increasing exposure to AAPA posts (Fig. 1). The first, a “Reduced Exposure” experiment, downranked all AAPA content, pushing it lower in their feeds. In contrast, the second, an “Increased Exposure” experiment, upranked attitudinally-aligned AAPA posts into participants’ feeds. Both experiments are randomized controlled trials. After providing informed consent and installing our web extension, participants were randomly assigned to one of the two experiments and then randomly divided into treatment or control conditions for each experiment. The overall experiment spanned ten days, with the first three days serving as a baseline period, during which no intervention was applied. During the subsequent seven days, participants in the treatment groups received the modified feed (downranked or upranked AAPA posts), while those in the control groups continued to receive the original, unmodified feed.
We assess outcomes at two different time points: post-experiment and in-feed. To measure post-experiment effects, we administered a survey asking participants how warm they feel towards people of the opposing party (0–100 degrees on a feeling thermometer) and about the emotions that they associate with reading social media over the previous week (5-point scale between “never” and “all the time”). We surveyed the participants both before and after the experiment, enabling us to quantify the impact of the intervention at the end of the week-long treatment (Sec. 5.2).
To measure in-feed effects, we added posts containing a small survey directly in the participant’s feed. These surveys, modeled on the ecological momentary assessment approach [45], resemble typical social media posts but are colored to stand out and contain a slider survey instrument. Each survey randomly selects one of two outcomes to assess: (1) affective polarization, where participants rate their feelings toward the opposing party, or (2) emotional responses, where participants report the extent to which they feel two emotions, one randomly selected positive emotion (excited or calm), and one randomly selected negative emotion (angry or sad). The in-feed surveys appeared with a decreasing probability after each feed intervention (post downranking or upranking) and were shown at an equivalent rate and position in the control conditions (Sec. 5.2).
2.3 Feed Reranking
To administer the interventions, we developed a browser extension for Google Chrome and Microsoft Edge that reranks participants’ feeds in real time [40]. The extension activates when participants access the “For You” feed on X. It intercepts the participants’ original feed and sends it to a remote backend, which scores and reranks the posts (Sec. B). This process adds a latency of about three seconds when opening the feed for the first time, which is consistent across all experimental conditions. To rerank the original feeds, we first identify political posts (Sec. B.3) and then determine whether those posts are classified as AAPA.
In the Reduced Exposure treatment condition, we downrank all AAPA posts in the participants’ feeds. If participants continue to scroll down far enough, e.g., across several incremental loads of more content, these demoted posts will appear (Fig. 9). Before downranking all the AAPA posts, we randomly select the position of one AAPA post and, if the in-feed survey is sampled to be included, insert the post with the survey in the next available position (Sec. A.2.1). This survey position is consistent across treatment and control conditions. Thus, participants in the control condition will see AAPA posts and the in-feed survey, while participants in the treatment condition will see the survey in the position where the first AAPA post used to be, but all AAPA posts are placed further down. The survey is limited to only a few times per day to avoid fatigue (Sec. 5.2). In the Increased Exposure experiment, since we lack access to the platform’s full post inventory, we sourced AAPA posts from the inventory of all posts in the “For You” feeds of all other study participants. We selected a random position in the original feed and upranked a random AAPA post to that position. The post containing the survey, if included, is inserted in the next available position. In the control condition, the survey is still added in the randomly sampled position, but no post is upranked. As in the Reduced Exposure experiment, the survey is placed consistently in the treatment and control conditions. During continuous scrolling sessions, this process repeats approximately every 30 posts (Sec. A.2.2).
The Reduced Exposure treatment downranked 85.4 AAPA posts on average (median = 22) per user per day, and the Increased Exposure upranked 10.9 AAPA posts on average (median = 5.43) per user per day.
3 Results
We recruited participants from two online survey platforms, CloudResearch and Bovitz-Forthright [46]. We targeted participants who use X, are 18 or older, reside in the United States, and self-identify as either Democrats or Republicans (Sec. C). After obtaining informed consent, participants installed our web extension. To ensure the participants had enough political content, we used a pre-screening task to filter for users whose “For You” feeds contained at least 5% of posts related to politics or social issues. The participants were recruited during the U.S. election season, between July 7th and August 14th, 2024, and instructed to use X daily only from the web interface.
We preregistered a target of at least 1,100 participants to complete the full study, broken down as 500 in the Increased Exposure experiment and 600 in the Reduced Exposure experiment based on power analyses performed on pilot studies. Toward this goal, a total of 1,662 participants completed the initial survey and qualified for inclusion in the study (Fig. 13). Participants were randomly assigned to either the Increased or Reduced Exposure group until the preregistered experiment-specific quotas were reached [47]. Of the qualified participants, N=1,256 completed the full study by responding to follow-up surveys the day after the end of the study: 727 in the Reduced Exposure experiment and 529 in the Increased Exposure experiment. Both treatment and control groups were well balanced across covariates, and there was no evidence of selective attrition between conditions (Sec. 5.3).
On average, participants viewed 158 posts per day (median = 94) and spent 43.8 minutes (median = 19.2 minutes) per day with a browser tab in the foreground on their X feed. During the baseline period, an average of 32% of the posts participants viewed were related to politics or social issues. Of these, 33.1% were classified as AAPA, which comprised an average of 10.6% of the total feed content.
The interventions had the intended effects (Fig. 11). During the baseline period, participants across all groups were exposed to political posts with similar average AAPA scores. However, after the intervention began, participants in the treatment condition of the Reduced Exposure experiment experienced a noticeable decrease in exposure to such posts (-91.2% compared to control), while those in the treatment condition of the Increased Exposure experiment were exposed to more AAPA posts (+23.1% compared to control), resulting in higher average scores.
Over the full ten days of the experiment, participants answered an average of 13.8 (median = 11) in-feed affective polarization questions in the Increased Exposure experiment and 12.6 (median = 9) in the Reduced Exposure experiment. Similarly, they answered an average of 14.0 (median = 11) emotion surveys in the Increased Exposure experiment and 12.5 (median = 9) in the Reduced Exposure experiment–each survey requiring ratings for two emotions.
In the post-survey, we asked participants whether the browser extension had impacted their experience on X. Across conditions, most participants reported that they did not notice any impact in their online experience (74.2%), while others guessed changes that were not part of the intervention (Sec. E.6). These results indicate a limited understanding of the experimental feed ranking intervention and suggest that it is less likely that demand effects might be influencing the outcomes.
Figure 2: Effects of reducing and increasing exposure to AAPA content in participants’ feeds on their feeling towards the out-party (feeling thermometer scale between 0: cold, and 100: warm) relative to the corresponding control group. Participants were surveyed within the feed during the intervention and off-platform after the experiment. The error bars represent 95% confidence intervals.
3.1 Affective polarization
Exposure to AAPA significantly influenced affective polarization (Fig. 2). Participants in the reduced AAPA exposure experiment reported a significant post-experiment effect, corresponding to an outparty warmth increase of 2.11 degrees (95% CIs: [0.15, 4.06]; P=0.035). Conversely, participants in the Increased Exposure experiment exhibited a symmetric reduction in affective warmth toward the opposing party of -2.48 degrees (95% CIs: [-4.79, -0.17]; P=0.036). We did not observe any significant heterogeneity in treatment effects across the preregistered moderators (Sec. E.5), including party identification, indicating that the effect was consistent and bipartisan. As is common [48], we contextualize these effects against the rising levels of partisan animosity in American society over time: our effects correspond to a decrease and increase in affective polarization equivalent to 3 and 3.6 years, respectively. We observed even stronger effects with the in-feed assessment, with Reduced Exposure causing an increase of 3.24 degrees (95% CIs: [1.21, 5.27]; P=0.002) and Increased Exposure causing a cooling of -2.56 degrees (95% CIs: [-4.53, -0.59]; P=0.011).
In an exploratory analysis, we investigated which AAPA factors were predictive of changes in affective polarization. Specifically, a higher proportion of posts supporting Partisan Animosity (95% CIs: [-13.49, -2.61]; P=0.003), Opposition to Bipartisan Cooperation (95% CIs: [-26.55, -4.40]; P=0.006), and Biased Evaluation of Politicized Facts (95% CIs: [-12.83, -2.21]; P=0.005;) predicted a colder feeling toward the opposing party by the end of the study (Fig. 17).
Figure 3: Effects of reducing and increasing exposure to AAPA content in participants’ feeds on their experiences of emotion (in-feed scale ranged from 0: “none at all” to 100: “extremely”, post-experiment scale ranged from 1: “never” to 5: “all the time”) relative to the corresponding control group. Participants were surveyed within the feed during the intervention and off-platform after the experiment. The filled circles represent statistical significance ($P<0.05$, adjusted for multiple hypothesis testing), and the error bars represent 95% confidence intervals.
3.2 Emotions
Changes to the feed algorithm caused in-feed, but not post-experiment, changes to participants’ negative emotions (Fig. 3). Reduced Exposure caused a change of -5.05 points (95% CIs: [-7.68, -2.42]; $P_{adj}$=0.002) for “angry” and -3.68 points (95% CIs: [-6.42, -0.94]; $P_{adj}$=0.024) for “sad.” Likewise, in the Increased Exposure, participants report a symmetric increase in negative emotions: 5.13 points (95% CIs: [2.05, 8.22]; $P_{adj}$=0.013) for “angry” and 4.38 points (95% CIs: [1.00, 7.76]; $P_{adj}$=0.036) for “sad.” Neither intervention had an effect on the positive emotions “calm” and “excited.” Post-experiment emotion effects were not significant, suggesting that the effects were likely short-lived.
3.3 Engagement
We tested whether our interventions impacted platform engagement (Sec. E.4). Neither experiment observed imbalances in attrition rates across conditions. Moreover, in both experiments, the treatment had no measurable impact on daily sessions—Increased Exposure: 1.65 control vs. 1.53 treatment, n.s.; Reduced Exposure: 1.53 vs. 1.49, n.s. Likewise, there was no significant effect on reshare rates—Increased Exposure: 0.52% vs. 1.07%, n.s.; decreased exposure: 0.74% vs. 0.46%, n.s. In addition, there was no effect on favorite rates—Increased Exposure: 4.48% vs. 5.46%, n.s; decreased exposure: 4.96% vs. 4.15%, n.s. We also observed that the differences between treatment and control groups in terms of the number of new posts, replies, and time spent on X during the experimental period were not statistically significant. These null results persisted when controlling for baseline behavior. While nominally consistently in the expected direction, none of the statistical tests were significant. It is possible that an even larger study could observe differences; however, our data suggests that the magnitude of any effect would be small.
In an exploratory analysis, we did observe that political content had higher engagement than non-political content, confirming the findings of previous work [49]. When focusing only on political content, we found that AAPA posts do not get more engagement than moderate political posts (Sec. E.4).
4 Discussion
In this study, we conducted a field experiment to investigate the effects of feed ranking algorithms that decrease or increase exposure to contact that polarizes. Our findings demonstrate that repeated exposure to AAPA posts, particularly during heightened political tensions such as months before the US election, significantly influences users’ affective polarization and emotional responses. A feed ranking algorithm that controls exposure to such content produces measurable changes, equivalent to mitigating approximately three years of affective polarization, even after just one week of treatment. Importantly, our results indicate that the intervention does not disproportionately impact any specific political leaning or demographic group, suggesting that the approach is bipartisan. The success of this method shows that it can be integrated into social media AIs to mitigate harmful societal consequences. We also present evidence that this impact can likely be achieved with a limited effect on business metrics like engagement.
Limitations of our study include a sample of partisan US users restricted to the web version of X during a specific period, which may affect the generalizability of the findings [50, 51]. Our population is limited to X users with at least 5% of political content (73.4% of the screened), but we intentionally selected this group to ensure the interventions can make a difference. Additionally, the experiment examines short-term effects—immediate in-feed responses and outcomes measured the day after the experiment—but does not address whether these effects persist over time. However, in scenarios where this ranking approach is integrated into the feed algorithm, the intervention would be continuously applied. Our reliance on in-feed surveys to assess immediate responses to stimuli follows established practices of ecological momentary assessment (EMA) [45]. To mitigate the risk of demand characteristics associated with the close proximity of the assessment, we followed best practices in designing the experiment [52], providing only essential information about the study’s purpose. Post-survey responses to an open-ended question about their experience during the study suggest that these precautions were effective, as participants did not report any impact on their experience attributable to their assigned condition. Describing the causal relation between social media and polarization as a unified phenomenon, of course, remains challenging: platforms are complex socio-technical systems, and small variations in design or exogenous factors can lead to different outcomes [53]. Future research should explore whether these effects persist over longer periods or in less politically charged contexts and examine the efficacy of feed reranking across different platforms and cultural settings.
As political polarization and societal division become increasingly linked to social media activity, our findings provide a potential pathway for platforms to address these challenges through algorithmic interventions. By embedding pro-democracy attitudes into AI-driven feeds [37, 39], we offer a concrete solution that could contribute to healthier democratic discourse, enhance social trust, and help mitigate the risks of deepening polarization.
Appendix A Interventions
A.1 Posts expressing Antidemocratic Attitudes and Partisan Animosity
We define AAPA based on eight variables identified [38] as potentially harmful to democratic functioning. While the previous study tested multiple interventions to measure the impact on these variables, we focus on content expressing support for these outcomes, following findings from an in-lab experiment [39].
The eight variables are:
  1. 1.
    Partisan animosity (v1): defined as dislike for opposing partisans [18].
  2. 2.
    Support for undemocratic practices (v2): defined as willingness to forgo democratic principles for partisan gain [61, 62]
  3. 3.
    Support for partisan violence (v3): defined as willingness to use violent tactics against outpartisans [63]
  4. 4.
    Support for undemocratic candidates (v4): defined as willingness to ignore democratic practices to elect inparty candidates [64]
  5. 5.
    Opposition to bipartisan cooperation (v5): defined as resistance to cross-partisan collaboration [65]
  6. 6.
    Social distrust (v6): defined as distrust of people in general [66]
  7. 7.
    Social distance (v7): defined as resistance to interpersonal contact with outpartisans [67]
  8. 8.
    Biased evaluation of politicized facts (v8): defined as skepticism of facts that favor the worldview of the other party [68]
These definitions, evaluated in previous work [39], were used to prompt the LLM (see Table 2). A post is classified as AAPA when at least four of these factors are present. Over the course of 10 days, we found that, on average, 30.6% of the political posts in the pre-intervention feed of participants were classified as AAPA (Fig. 4). The distribution of AAPA varied by participants’ political orientation, with 34.25% of political posts in Republican feeds and 28.54% in Democratic feeds classified as AAPA (Fig. 5). The most common factors are Social Distance, Biased Evaluation of Politicized Facts, and Partisan Animosity (Fig. 6).
Fig. 7 shows the correlation between these factors, showing that some of them have a high chance of occurring together. Fig. 8 shows that these factors cluster in 3 groups with {v2, v3, v4}, {v5}, and {v1, v6, v7, v8} tendency to appear together.
Figure 4: Left: Distribution of the AAPA scores assigned to political posts in the “For you” feed on X (pre-intervention). The green bars represent the fraction of non-AAPA, while the orange bars describe the AAPA posts recommended by X.
Figure 5: AAPA score distribution for political posts, by party.
Figure 6: Distribution of AAPA factors in political posts seen by participants.
Figure 7: Correlation between the eight AAPA variables based on the co-occurrence in all the posts recommended by X to all participants in their ”For You” feed.
Figure 8: PCA analysis of the eight AAPA variables. Reduction of the binary vectors representing the factors present in all the posts recommended by X to all participants in their ”For You” feed. The eight factors cluster into three groups.
A.2 Feed reranking implementation
We developed a browser extension to run the field experiment intervening in real-time in the participants’ original feeds [47]. Our extension was distributed through the Google Chrome Store and supported Chrome and Edge browsers. Once the participants consented to participate, the extension is enabled every time they visit their feed on X. The extension is enabled only when the participants use the “For You” feed, which is the feed generated by the recommendation system. Participants can read their chronological feed at any time, but a tooltip regularly reminds them to switch to the “For You” tab to participate in the study. Our implementation relies on the browser extension to intercept the feed and on a remote backend to rescore and rerank the posts.
Every time the user loads the feed, the extension receives a working set of 35 posts (5 of which are ads) from the server, and for each of these batches, the reranking procedure works as follows. First, the extension intercepts the response returned from the server and extracts the text of all the posts. If the post contains a link, we append the title and description visible in the preview with the prefix “Attached article’s description: ” and discard the URL. If the text quotes another post, we append the quoted text with the prefix “Quoting: .” Second, since our reranking logic is applied only to political posts, our reranking service selects all relevant posts in the feed using a RoBERTa-based political classifier (Sec. B.3). Filtering for political posts with a lighter model allows us to reduce the latency by scoring only the relevant posts. Then, the political posts are scored in real-time using GPT-3.5 using prompts based on previous work [39] and summarised in Table 2. For each post, the resulting data is a vector of 8 binary indicators flagging the presence of each factor. The final score of each post can vary between 0 and 8, and it is obtained by counting how many factors the classifier identified.
Finally, depending on the participant’s exposure group, we apply the relevant intervention and return the modified feed to the browser, potentially with the in-feed survey. Each in-feed survey contained a slider for participants to respond using a 0–100 scale on the feeling thermometer. The design required two control conditions since the Reduced Exposure intervention placed the in-feed survey instrument where downranked content had previously been, and the Increased Exposure intervention placed the in-feed assessment instrument where upranked content would have been inserted.
A.2.1 Reduced Exposure feed
Participants in the treatment condition of the Reduced Exposure group receive a feed with less AAPA content. This is achieved by penalizing these posts and moving them to a lower position in the feed. Once the relevant posts have been scored, we proceed as follows. First, we select one random AAPA item and flag the next position as the potential location for the in-feed survey. Then, if the participant is currently in the treatment condition, we downrank all the problematic posts. If the participant is in the control condition, we return the feed with the original ranking. The penalization offset for downranked posts depends on their score (position × score × 10) to ensure that posts with higher scores require more effort to become visible in the feed. Our implementation does not remove the posts, and the participants can read the downranked content by scrolling down the feed in a continuous session. This implementation ensures that the selection logic for the survey position is equivalent for the two experimental conditions, differing only in the downranking of the problematic posts for the treatment group. Finally, the in-feed survey is injected in the assigned position if the current feed load is selected for inserting the survey. If the current feed does not contain any problematic posts, there is no intervention, and we do not add any surveys. The effect of the intervention is visible in Fig. 9.
A.2.2 Increased Exposure feed
The participants in the Increased Exposure group are presented with a treatment opposite to the Reduced Exposure group, as they can find AAPA content in the top positions of their feed. The treatment condition of this group receives a feed where AAPA posts are upranked. Since we cannot access the full inventory available on the platform to generate the custom recommendation, we source the AAPA posts from a custom inventory compiled from all the items recommended to the participants enrolled in the study. In particular, to ensure recency and relevance, we limit the candidate set to posts not already seen by the current participant and recommended by the feed algorithm to other participants with the same political leaning in the last 24 hours. This list may include posts recommended to the same participants that were not previously displayed. The content to uprank is randomly selected from the 100 posts with the highest AAPA score and matching these characteristics. Once the post is selected, we select one random position in the feed, and if the participant is in the treatment condition, the item will be inserted. Alternatively, if the participant is assigned to the control condition, the process is interrupted, and the post is not added. Despite not adding the post, mirroring the same procedure for the control group ensures the similarity of the two conditions in all respects, including potential server latency. Finally, if the event is sampled to include an in-feed survey, we append the in-feed survey in the next available position. The effect of the intervention is visible in Fig. 10.
Figure 9: Reduced Exposure: Distribution of the average number of political posts that each participant saw by day for each score. During the treatment period, participants experienced a significant reduction (-91.2%) in their exposure to AAPA content ($\geq$ 4 factors). While AAPA content was downranked, participants who scrolled far enough through their feeds still encountered a small amount of AAPA posts.
Figure 10: Increased Exposure: Distribution of the average number of political posts that each participant saw by day for each score. During the treatment period, participants experienced a significant increase (+23.1%) in their exposure to AAPA content ($\geq$ 4 factors).
Figure 11: Average AAPA score for political posts by day and condition; first averaged per participant and then averaged across all users in the corresponding group. The results demonstrate that the intervention increased and decreased AAPA exposure, as intended.
Figure 12: Distribution of the number of AAPA posts upranked and downranked per participant, per day, by in the feeds of participants assigned to the corresponding treatment groups. The x-axis is on a logarithmic scale. The Reduce Exposure treatment downranked approximately an order of magnitude more AAPA posts per participant than the number of posts the Increased Exposure treatment upranked.
Appendix B Experimental platform
We developed a custom experimental platform focused on performance to ensure the experiment introduces minimal disruption to the user experience on X. Our implementation consists of a browser extension and a REST backend deployed as a cloud service.
B.1 Browser extension
The browser extension handles editing the X feed and managing all interactions with the participants, including the onboarding steps. It is compatible with Google Chrome and Microsoft Edge and requires only storage permissions to save the user identifier.
The extension manages the ranking manipulation by modifying the content available on the X feed in real-time. Our implementation allows us to inject custom code on the main page to override the methods of the native XMLHttpRequest object. This approach allows us to intercept the communication with the server and modify the execution flow of the requests. Our browser extension subscribes to the requests endpoints used to retrieve the new content to add to the feed, interrupt the execution flow, send the server response to our backend for scoring and reranking, and finally restore the execution flow by calling the original methods with the response customized by our server. This approach requires careful understanding of the communication protocol with the server, but it offers significant advantages: first, the content can be modified by editing structured JSON objects without manipulating the final HTML, and second, by selecting what requests to intercept, it allows logging the interactions with the web page tracked by X without adding custom listeners.
Additionally, the browser extension interacts with the main page by editing the DOM to add visual elements. These widgets include the in-feed surveys that are added to the feed as regular posts and a top fixed banner used to notify the participant about possible registration errors or that the experiment is completed, with the invitation to take the final survey and remove the extension.
Finally, the extension manages the onboarding flow to ensure that participants can register and join the study only by following links that automatically guide them in installing the extension and setting the identifier from the recruitment platform [40].
B.2 Backend
The backend is responsible for reranking the feed, logging the client events, sending daily participation reminders by email, and orchestrating the participants’ registration. The server is written in Python and Java using MongoDB and Memcached for storage and caching. When the browser extension interrupts the execution flow to request the reranking of the feed, X’s rotating circular loading indicator is extended until our reranking is complete. The implementation choices are guided by minimizing the response time to ensure minimal destruction of the user experience.
The feed reranking proceeds in four steps: first, the response from X’s server, typically containing 35 posts, is sent to our backend and parsed to extract the text from the posts. The implementations focused on individual posts and in-feed conversations, ignoring the other types of content, such as ads or users’ recommendations. Second, all the posts are evaluated by a political classifier (Sec. B.3) to filter the posts that must be scored for the presence of the eight factors problematic for democratic functioning. This service runs as a standalone Python web service, accelerated by GPU, that introduces, on average, a latency of 490 ms (Std = 111 ms). Then, the political posts are dispatched to a Java server that manages an HTTP connection pool with OpenAI APIs. The service batches the posts to reduce the number of requests and uses a pool of tokens to minimize latency and avoid the rate limit exceptions. Finally, the scores are parsed and returned to the main backend, which applies the downranking intervention and adds the in-feed surveys. These steps are replicated in the same way in the control condition, but in that case, the backend returns the original feed. The downranked posts are saved in a cache (Memcached) and added to the feed when the user reaches the new position in a continuous scrolling session.
The backend also stores the logs of interactions with the feed. When a participant interacts with the feed on X, the extension regularly shares with the backend what actions are performed on the main page. This data is limited to the actions on the main feed and does not include any private content. The actions recorded include when a post enters the main viewport, likes, favorites, reposts, clicks on external links, and the time spent on the browser tab.
To facilitate communication with the participants in the registration flow, they are invited to share an email address. The platform uses this address to send automatic messages that guide the participant in completing the study successfully. These messages include a copy of the onboarding instructions and the consent form for future reference, a daily reminder sent at 6 PM (participant’s local time) inviting the participant to check their X feed, and the end-of-study notification with the link to the post-survey and the instruction to remove the browser extension.
Finally, the backend orchestrates the registration flow of the participants [47]. Since the browser extension does not support passing custom installation parameters, the backend is responsible for guiding the onboarding flow to configure the extension for the current participant. After the termination of the pre-survey, the participants are redirected to our web server using a URL that encodes all the parameters relevant for the extension to work, such as the identifier and political position. Our server adds a browser cookie with this data and displays the step-by-step instructions for installing the extension. Participants are then redirected to the Chrome store, where they install the browser extension. As soon as the extension is installed, it opens a browser tab and loads to a welcome page hosted on our coordination server, where it has access to the cookies information previously set. This approach allows the extension to save these parameters in persistent storage without forcing the participants to enter this information manually after the installation of the extension. Lastly, the welcome message contains a link to redirect the participants back to the recruitment platform to set the task as completed.
B.3 Political Classifier
Since our downranking intervention applies only to political posts, we developed a classifier to prefilter the relevant posts. Since GPT has a pay-per-use business model and calling an external API service call can introduce additional latency, we ensure that GPT is used to score only the content in which the eight variables are relevant. The political classifier is a distilled version of GPT-4 by fine-tuning a RoBERTa model to identify political posts, including mentions of officials and activists, social issues, news, and current events.
The posts used to train the model were collected with the Twitter API in July 2023. To select the posts, we replicated the reverse chronological feeds of users across the political spectrum for two weeks and took a random sample of 50,000 posts. We first selected 300 users and then collected the posts by the users they followed.
To select the users, we first obtained the followers of all members of Congress. Then, we categorized each user in the superset of followers based on how many Republican and Democrat members of Congress they followed: (a) far-left: followed 4-6 Democrats and no Republicans; (b) left: followed 1-3 Democrats and no Republicans; (c) middle: followed 1-3 Democrats and 1-3 Republicans; (d) far-middle: followed 4-6 Democrats and 4-6 Republicans; (e) right: follow no Democrats and 1-3 Republicans; (f) far-right: followed no Democrats and 4-6 Republicans. We started with members of Congress since their political affiliations are known and are often used to estimate Twitter users’ political leaning [69, 70]. Next, we removed follower accounts that are likely to be bots. We used the Botometer API and filtered out accounts with Complete Automation Probability above 0.6 [71]. Finally, we took a stratified sample of 300 accounts, selecting 50 users from each of the six groups defined above.
Next, for each of the selected users, we collected all of the accounts they followed and collected all the posts by those users as far back as we could. Then, to replicate the reverse chronological feed of each user, we merged all posts by the accounts they followed and sorted from most to least recent. We retain only the English posts that would have appeared in their reverse chronological feeds in the previous two weeks. Finally, we merged the two-week feeds of all users, removed duplicate posts, and took a random sample of 50,000 posts. We later randomly split this dataset into 45,000 posts for training and 5,000 posts for testing.
We annotated all 50,000 posts by prompting GPT-4 using a definition proposed by the Pew Research Center [44]:
Political content on Twitter is varied and can be about officials and
activists, social issues, or news and current events.
Looking at the following tweet, would you categorize it as POLITICAL
or NOT POLITICAL content?
Answer 1 if it is POLITICAL, 0 otherwise.
This step annotated 34.2% of the posts as political and 65.8% as non-political. We used this dataset to fine-tune a pre-trained RoBERTa model (roberta-base).
We evaluate the model on a held-out dataset composed of 5,000 posts annotated by GPT-4 and a random subset of 300 posts annotated by humans. For each post, we collect three labels using Prolific and assign the value based on majority voting. The model evaluated on these testing sets achieves an F1 score of 93.6% on the GPT-4 labels and an F1 score of 91.3% on the human ground truth.
Ground truth Test set size Precision [95% CI] Recall [95% CI] F1 score
GPT-4 5,000 0.937 [0.931, 0.942] 0.920 [0.911, 0.928] 0.936
Human labels 300 0.921 [0.868, 0.966] 0.905 [0.847, 0.954] 0.913
Table 1: Performance of the fine-tuned RoBERTa political classifier. Bootstrapped 95% confidence intervals.
B.4 Scoring service
The scoring service is a Java backend responsible for returning the number of each political post’s eight problematic factors. It is designed to minimize latency. When a new request is received, it proceeds as follows: after the political classifier identifies the posts to score, they are sent to this service in one request as a list of textual items. This service first splits the list into chunks of at most ten items and then compiles eight prompts for each of them. The prompts are in the format:
Do the following messages express partisan animosity?
Partisan animosity is defined as "dislike for opposing partisans".
FORMAT:
The input messages are given as JSON lines in the format
{"id": <message_id>, "message": <message>}.
The output must be a JSON array of objects in the format
[{"id": <message_id>, "answer": <YES or NO>}, ... ].
INPUT MESSAGES:
The prompt starts with the question and the definition of the factor as described in Table 1, specifies the input and output format, and ends with the list of posts encoded in JSON objects, one for each line. The posts are grouped with at most ten items per prompt as a tradeoff between making too many API calls and keeping the response time low. This implementation ensures that a feed containing ten political posts results in 8 requests to OpenAI instead of 80.
These requests run in parallel using a custom connection pool that shares HTTP connections and iterates round-robin on multiple API tokens. Each request takes a few seconds to receive a response, but each connection has a maximum waiting time of 8 seconds to ensure a slow API call does not disrupt the user experience. When the timeout is reached, the event is logged, and the factor is set as non-present. To ensure a fast response from OpenAI, we use GPT-3.5 Turbo, which offers a good tradeoff between speed and accuracy. We use a fixed seed for reproducibility purposes.
Since many popular posts appear in multiple feeds, to avoid scoring multiple times the same text, each post score is cached in a backend database and reused when necessary without including the post in the request to OpenAI.
Variable Definition
Partisan Animosity Do the following messages express partisan animosity? Partisan animosity is defined as ”dislike for opposing partisans”.
Support for Undemocratic Practices Do the following messages express support for undemocratic practices? Support for undemocratic practices is defined as ”willingness to forgo democratic principles for partisan gain”. Undemocratic practices are undemocratic tendencies or actions such as reducing polling stations in areas that support their opponents, attacking the independence of the judiciary, undermining the free press, challenging the legitimacy of election results, or encouraging political violence.
Support for Partisan Violence Do the following messages express support for partisan violence? Support for partisan violence is defined as a ”willingness to use violent tactics against outpartisans”. Examples of partisan violence include sending threatening and intimidating messages to the opponent party, harassing the opponent party on the Internet, using violence in advancing their political goals or winning more races in the next election.
Support for Undemocratic Candidates Do the following messages express support for undemocratic candidates? Support for undemocratic candidates is defined as ”willingness to ignore democratic practices to elect inparty candidates.” Undemocratic candidates often support undemocratic practices such as reducing polling stations in areas that support their opponents, attacking the independence of the judiciary, undermining the free press, challenging the legitimacy of election results, or encouraging political violence.
Opposition to Bipartisan Cooperation Do the following messages express opposition to bipartisanship? Opposition to bipartisanship is defined as ”resistance to cross-partisan collaboration”.
Social Distrust Do the following messages express social distrust? Social distrust is defined as ”distrust of people in general”.
Social Distance Do the following messages express social distance? Social distance is defined as ”resistance to interpersonal contact with outpartisans”. Messages that increase social distance may contain terms that increase distrust, distance, insecurity, hate, prejudice, or discrimination.
Biased Evaluation of Politicized Facts Do the following messages express a biased evaluation of politicized facts? Biased evaluation of politicized facts is defined as ”skepticism of facts that favor the worldview of the other party”. Messages supporting a biased evaluation of politicized facts may partially present political facts or discuss a controversial issue with a certain political stance.
Table 2: Definition of the eight factors used in the GPT prompts.
Appendix C Participant characteristics
Figure 13: Summary of the participants’ recruitment flow.
To enroll in the experiment, participants had to take a screening task by installing a browser extension. The extension opened X, asked participants to log in if necessary, scrolled through their feed automatically, and verified that at least 5% of the posts in their “For You” feed were about politics or social issues. Fig. 15 summarizes the distribution of users by the fraction of political content. Around 73% of the participants qualified and were invited to the full study.
Figure 14 shows the distributions of some of the participants’ demographic information collected in the pre-survey, including race, education level, perceived socioeconomic status, and income.
Figure 14: Distribution of the participants’ demographic characteristics.
Figure 15: Distribution of the fraction of political content in the feed of the participants screened. The green area represents the participants who were qualified to join the study.
C.1 Covariate Balance
Tables 3 and 4 show univariate analyses of covariate balance in the Reduced and Increased Exposure experiments, respectively. Omnibus test of balance using randomization inference are reported in Section 5. We do not observe any evidence of imbalance.
Treatment Control Difference p(diff != 0) p adjusted
Party: Democrat (%) 0.672 0.692 -0.020 0.590 1.000
Party: Republican (%) 0.328 0.308 0.020 0.590 1.000
Party strength: Weak (%) 0.497 0.470 0.027 0.507 1.000
Party strength: strong (%) 0.503 0.530 -0.027 0.507 1.000
Party importance (0 - 100) 64.747 68.430 -3.684 0.100 1.000
Gender: Female (%) 0.459 0.467 -0.008 0.851 1.000
Gender: Male (%) 0.522 0.513 0.009 0.830 1.000
Gender: Other (%) 0.019 0.020 -0.001 0.920 1.000
Age 41.081 42.434 -1.353 0.223 1.000
Race: white (%) 0.784 0.742 0.043 0.212 1.000
Race: Black (%) 0.153 0.149 0.004 0.886 1.000
Race: Hispanic (%) 0.094 0.109 -0.016 0.523 1.000
Race: Asian (%) 0.078 0.083 -0.005 0.831 1.000
Race: Other (%) 0.038 0.030 0.008 0.595 1.000
Education: High school or less (%) 0.144 0.129 0.015 0.596 1.000
Education: Some college (%) 0.284 0.318 -0.034 0.364 1.000
Education: Bachelor’s degree (%) 0.356 0.348 0.009 0.823 1.000
Education: Master’s, PhD, JD, MD (%) 0.216 0.205 0.010 0.753 1.000
Income: less than \$30,000 (%) 0.125 0.166 -0.041 0.153 1.000
Income: \$30,000 - \$80,000 (%) 0.394 0.437 -0.043 0.274 1.000
Income: \$80,000 - \$125,000 (%) 0.303 0.222 0.081 0.021 1.000
Income: more than \$125,000 (%) 0.178 0.175 0.003 0.932 1.000
SES Ladder: Bottom rungs (%) 0.166 0.182 -0.016 0.588 1.000
SES Ladder: Middle rungs (%) 0.741 0.725 0.015 0.664 1.000
SES Ladder: Top rungs (%) 0.094 0.093 0.001 0.965 1.000
Recruitment platform: Bovitz (%) 0.659 0.599 0.060 0.122 1.000
Recruitment platform: CloudResearch (%) 0.341 0.401 -0.060 0.122 1.000
Table 3: Univariate Covariate Balance: Reduced Exposure Experiment
Treatment Control Difference p(diff != 0) p adjusted
Party: Democrat (%) 0.710 0.622 0.088 0.043 0.289
Party: Republican (%) 0.290 0.378 -0.088 0.043 0.289
Party strength: Weak (%) 0.454 0.483 -0.029 0.533 1.000
Party strength: strong (%) 0.546 0.517 0.029 0.533 1.000
Party importance (0 - 100) 67.429 63.935 3.494 0.158 0.501
Gender: Female (%) 0.542 0.448 0.094 0.042 0.289
Gender: Male (%) 0.441 0.526 -0.085 0.066 0.289
Gender: Other (%) 0.017 0.026 -0.009 0.490 1.000
Age 42.916 41.470 1.446 0.266 0.918
Race: white (%) 0.777 0.757 0.021 0.596 1.000
Race: Black (%) 0.134 0.139 -0.005 0.883 1.000
Race: Hispanic (%) 0.080 0.135 -0.055 0.056 0.289
Race: Asian (%) 0.092 0.048 0.045 0.058 0.289
Race: Other (%) 0.034 0.043 -0.010 0.581 1.000
Education: High school or less (%) 0.139 0.139 0.000 0.988 1.000
Education: Some college (%) 0.361 0.383 -0.021 0.635 1.000
Education: Bachelor’s degree (%) 0.319 0.322 -0.002 0.956 1.000
Education: Master’s, PhD, JD, MD (%) 0.181 0.157 0.024 0.486 1.000
Income: less than \$30,000 (%) 0.193 0.165 0.028 0.430 1.000
Income: \$30,000 - \$80,000 (%) 0.412 0.422 -0.010 0.827 1.000
Income: \$80,000 - \$125,000 (%) 0.164 0.252 -0.088 0.019 0.289
Income: more than \$125,000 (%) 0.231 0.161 0.070 0.056 0.289
SES Ladder: Bottom rungs (%) 0.231 0.217 0.014 0.723 1.000
SES Ladder: Middle rungs (%) 0.685 0.722 -0.037 0.384 1.000
SES Ladder: Top rungs (%) 0.084 0.061 0.023 0.334 1.000
Recruitment platform: Bovitz (%) 0.576 0.578 -0.003 0.954 1.000
Recruitment platform: CloudResearch (%) 0.424 0.422 0.003 0.954 1.000
Table 4: Univariate Covariate Balance: Increased Exposure Experiment
Appendix D Adjustment for multiple comparisons
For the two groups (experiments), we independently implement an FDR adjustment [72], grouping the variables as follows:
  • $K_{1}$ primary outcomes: Not adjusted,
  • $K_{2}$ secondary outcomes: Sharpened FDR-adjusted p-values with $K_{1}+K_{2}$ outcomes,
  • $K_{3}$ tertiary outcomes: Sharpened FDR-adjusted p-values with $K_{1}+K_{2}+K_{3}$ outcomes.
We also use correct for multiple comparisons when examining heterogeneous treatment effects.
  • Primary tests of heterogeneity: We used sharpened FDR-adjusted p-values with $L_{x}*K_{x}$ hypothesis tests, where $L_{x}$ is the number of primary moderators for the corresponding group, and $K_{x}$ is the number of outcomes in the corresponding group (i.e., $K_{1}$, $K_{2}$, or $K_{3}$ above).
  • Secondary tests of heterogeneity: We specify the secondary moderators that we considered for each group of outcomes (primary, secondary, tertiary). We used sharpened FDR-adjusted p-values with $(L_{x}+L_{y})*K_{x}$ hypothesis tests, where $L_{x}$ is the number of primary and $L_{y}$ is the number of secondary moderators for the corresponding group, and $K_{x}$ is the number of outcomes in the corresponding group (i.e., $K_{1}$, $K_{2}$, or $K_{3}$ above).
  • Auxiliary tests: Not adjusted, generally these are variables for which statistical significance should not matter.
Appendix E Supplementary Results
E.1 Affective Polarization: Regression Tables
The preregistered regression analysis for the in-feed effect is reported in Table 5, and the regression analyses for the post-experiment effect are reported in Table 6.
Table 7 and Table 8 show the equivalent regressions but include the entire population (N=1,256), regardless of whether the participants completed only the post-survey or only at least one in-feed survey. For both experiments, the treatment effect remains statistically significant. We note that in this sample there is also no evidence of covariate imbalance (reduce, post-experiment: $p=0.86$; reduce, in-feed: $p=0.59$; increase, post-experiment: $p=0.77$; increase, in-feed: $p=0.18$), asymmetric attrition rates (reduce, post-experiment: $p=0.55$; reduce, in-feed: $p=0.28$; increase, post-experiment: $p=0.4$; increase, in-feed: $p=0.56$), or asymmetric attrition patterns (reduce, post-experiment: $p=0.49$; reduce, in-feed: $p=0.69$; increase, post-experiment: $p=0.78$; increase, in-feed: $p=0.98$). Testing procedures are described in Section 5.3.
In-Feed
Reduced Exposure Increased Exposure
(Intercept) $0.90$ $3.06^{*}$
$[-1.08;2.87]$ $[1.08;5.05]$
Condition: Treatment $3.24^{*}$ $-2.56^{*}$
$[1.21;5.27]$ $[-4.53;-0.59]$
Baseline: Mean Outparty Feeling Thermometer $0.87^{*}$ $0.84^{*}$
$[0.83;0.92]$ $[0.79;0.88]$
Recruitment Platform: CloudResearch $-0.58$ $-1.58$
$[-2.69;1.54]$ $[-3.58;0.43]$
Num. obs. $6116$ $4778$
Num. groups. $622$ $468$
Null hypothesis value outside the confidence interval.
Table 5: Affective Polarization: In-Feed. Regression tables of the analysis of the affective polarization outcome for the in-feed survey responses.
Post-Experiment
Reduced Exposure Increased Exposure
(Intercept) $1.26$ $3.89^{*}$
$[-0.76;3.27]$ $[1.47;6.31]$
Condition: Treatment $2.11^{*}$ $-2.48^{*}$
$[0.15;4.06]$ $[-4.79;-0.17]$
Pre-survey: Outparty Feeling Thermometer $0.81^{*}$ $0.83^{*}$
$[0.77;0.85]$ $[0.77;0.88]$
Recruitment Platform: CloudResearch $-0.76$ $-2.72^{*}$
$[-2.79;1.28]$ $[-5.06;-0.37]$
Num. obs. $622$ $468$
Null hypothesis value outside the confidence interval.
Table 6: Affective Polarization: Post-Experiment. Regression tables of the analysis of the affective polarization outcome for the post-experiment survey responses.
In-Feed
Reduced Exposure Increased Exposure
(Intercept) $0.81$ $2.42^{*}$
$[-1.13;2.75]$ $[0.44;4.39]$
Condition: Treatment $3.30^{*}$ $-2.45^{*}$
$[1.30;5.30]$ $[-4.42;-0.48]$
Baseline: Mean Outparty Feeling Thermometer $0.88^{*}$ $0.85^{*}$
$[0.83;0.92]$ $[0.81;0.89]$
Recruitment Platform: CloudResearch $-0.58$ $-0.82$
$[-2.68;1.51]$ $[-2.84;1.19]$
Num. obs. $6273$ $4871$
Num. groups. $649$ $499$
Null hypothesis value outside the confidence interval.
Table 7: Full population – Affective Polarization: In-Feed. Regression tables of the analysis of the affective polarization outcome for the in-feed survey responses.
Post-Experiment
Reduced Exposure Increased Exposure
(Intercept) $1.72$ $3.63^{*}$
$[-0.14;3.59]$ $[1.34;5.91]$
Condition: Treatment $1.87^{*}$ $-2.24^{*}$
$[0.05;3.69]$ $[-4.40;-0.08]$
Pre-survey: Outparty Feeling Thermometer $0.81^{*}$ $0.84^{*}$
$[0.77;0.85]$ $[0.80;0.89]$
Recruitment Platform: CloudResearch $-0.79$ $-3.07^{*}$
$[-2.72;1.14]$ $[-5.30;-0.84]$
Num. obs. $727$ $529$
Null hypothesis value outside the confidence interval.
Table 8: Full population – Affective Polarization: Post-Experiment. Regression tables of the analysis of the affective polarization outcome for the post-experiment survey responses.
E.2 Emotions: Regression Tables
The preregistered regression analysis for the in-feed effect in the reduced exposure is reported in Table 9, while for the increased exposure is summarized in Table 10. The post-experiment effect for the reduced exposure experiment is summarized in Table 11 and for the increased exposure in Table 12.
Table 13, Table 14, Table 15, and Table 16 report the equivalent regressions but include the entire population (N=1,256), regardless of whether the participants completed only the post-survey or only at least one in-feed survey. The results are qualitatively similar, the variation in negative emotions is statistically significant in both experiments: increased in the increased exposure experiment and reduced in the reduced exposure experiment.
Reduced Exposure, In-Feed
Angry Sad Excited Calm
(Intercept) $9.26^{*}$ $7.77^{*}$ $12.53^{*}$ $17.67^{*}$
$[6.67;11.85]$ $[5.04;10.49]$ $[9.03;16.04]$ $[12.78;22.57]$
Condition: Treatment $-5.05^{*}$ $-3.68^{*}$ $1.47$ $2.92$
$[-7.68;-2.42]$ $[-6.42;-0.94]$ $[-1.91;4.84]$ $[-0.15;5.99]$
Baseline: Mean Answer $0.54^{*}$ $0.68^{*}$ $0.75^{*}$ $0.68^{*}$
$[0.47;0.62]$ $[0.60;0.75]$ $[0.68;0.82]$ $[0.61;0.75]$
Recruitment Platform: CloudResearch $0.56$ $2.42$ $-5.05^{*}$ $-3.09$
$[-2.16;3.27]$ $[-0.41;5.24]$ $[-8.54;-1.55]$ $[-6.25;0.08]$
Num. obs. $3057$ $3007$ $2990$ $3098$
Num. groups. $574$ $566$ $572$ $567$
Null hypothesis value outside the confidence interval.
Table 9: Emotions: Reduced Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the reduced exposure experiment.
Increased Exposure, In-Feed
Angry Sad Excited Calm
(Intercept) $5.98^{*}$ $6.43^{*}$ $14.91^{*}$ $20.87^{*}$
$[3.05;8.90]$ $[3.07;9.79]$ $[10.94;18.89]$ $[15.33;26.40]$
Condition: Treatment $5.13^{*}$ $4.38^{*}$ $-0.09$ $-2.69$
$[2.05;8.22]$ $[1.00;7.76]$ $[-3.75;3.57]$ $[-6.09;0.70]$
Baseline: Mean Answer $0.56^{*}$ $0.55^{*}$ $0.68^{*}$ $0.66^{*}$
$[0.47;0.64]$ $[0.46;0.64]$ $[0.60;0.76]$ $[0.59;0.74]$
Recruitment Platform: CloudResearch $1.46$ $5.10^{*}$ $-6.31^{*}$ $-3.93^{*}$
$[-1.65;4.57]$ $[1.70;8.50]$ $[-9.99;-2.62]$ $[-7.35;-0.51]$
Num. obs. $2451$ $2380$ $2434$ $2409$
Num. groups. $443$ $442$ $441$ $453$
Null hypothesis value outside the confidence interval.
Table 10: Emotions: Increased Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the increased exposure experiment.
Reduced Exposure, Post-Experiment
Angry Sad Excited Calm
(Intercept) $1.20^{*}$ $1.06^{*}$ $1.28^{*}$ $2.12^{*}$
$[1.01;1.39]$ $[0.88;1.24]$ $[1.08;1.48]$ $[1.89;2.36]$
Condition: Treatment $-0.12$ $0.02$ $0.04$ $0.03$
$[-0.26;0.02]$ $[-0.11;0.14]$ $[-0.11;0.18]$ $[-0.12;0.18]$
Pre-survey Answer $0.46^{*}$ $0.45^{*}$ $0.51^{*}$ $0.40^{*}$
$[0.39;0.53]$ $[0.39;0.52]$ $[0.44;0.57]$ $[0.34;0.47]$
Recruitment Platform: CloudResearch $0.06$ $0.05$ $-0.07$ $-0.04$
$[-0.08;0.21]$ $[-0.08;0.18]$ $[-0.22;0.08]$ $[-0.20;0.11]$
Num. obs. $617$ $617$ $617$ $617$
Null hypothesis value outside the confidence interval.
Table 11: Emotions: Reduced Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses in the reduced exposure experiment.
Increased Exposure, Post-Experiment
Angry Sad Excited Calm
(Intercept) $0.88^{*}$ $0.95^{*}$ $1.28^{*}$ $2.14^{*}$
$[0.67;1.10]$ $[0.72;1.17]$ $[1.05;1.51]$ $[1.86;2.42]$
Condition: Treatment $0.01$ $0.06$ $0.06$ $-0.12$
$[-0.15;0.17]$ $[-0.10;0.22]$ $[-0.10;0.22]$ $[-0.29;0.05]$
Pre-survey Answer $0.59^{*}$ $0.54^{*}$ $0.53^{*}$ $0.41^{*}$
$[0.52;0.67]$ $[0.46;0.62]$ $[0.45;0.61]$ $[0.33;0.48]$
Recruitment Platform: CloudResearch $0.08$ $0.17^{*}$ $-0.25^{*}$ $-0.12$
$[-0.08;0.24]$ $[0.01;0.34]$ $[-0.42;-0.08]$ $[-0.28;0.05]$
Num. obs. $472$ $472$ $472$ $472$
Null hypothesis value outside the confidence interval.
Table 12: Emotions: Increased Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses in the increased exposure experiment.
Reduced Exposure, In-Feed
Angry Sad Excited Calm
(Intercept) $9.14^{*}$ $7.44^{*}$ $12.11^{*}$ $17.23^{*}$
$[6.59;11.69]$ $[4.79;10.09]$ $[8.65;15.56]$ $[12.36;22.11]$
Condition: Treatment $-4.65^{*}$ $-3.28^{*}$ $1.39$ $2.56$
$[-7.24;-2.06]$ $[-5.96;-0.60]$ $[-1.94;4.73]$ $[-0.47;5.60]$
Baseline: Mean Answer $0.54^{*}$ $0.68^{*}$ $0.74^{*}$ $0.69^{*}$
$[0.47;0.61]$ $[0.61;0.76]$ $[0.67;0.82]$ $[0.63;0.76]$
Recruitment Platform: CloudResearch $0.62$ $2.52$ $-4.23^{*}$ $-3.10$
$[-2.07;3.32]$ $[-0.26;5.30]$ $[-7.71;-0.76]$ $[-6.24;0.04]$
Num. obs. $3120$ $3081$ $3062$ $3162$
Num. groups. $595$ $587$ $596$ $587$
Null hypothesis value outside the confidence interval.
Table 13: Full population – Emotions: Reduced Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the reduced exposure experiment.
Increased Exposure, In-Feed
Angry Sad Excited Calm
(Intercept) $6.03^{*}$ $6.21^{*}$ $14.45^{*}$ $20.12^{*}$
$[3.16;8.91]$ $[2.95;9.47]$ $[10.53;18.37]$ $[14.72;25.52]$
Condition: Treatment $4.89^{*}$ $4.61^{*}$ $0.23$ $-2.76$
$[1.86;7.91]$ $[1.30;7.92]$ $[-3.37;3.83]$ $[-6.08;0.57]$
Baseline: Mean Answer $0.55^{*}$ $0.57^{*}$ $0.68^{*}$ $0.67^{*}$
$[0.47;0.63]$ $[0.49;0.66]$ $[0.60;0.76]$ $[0.60;0.75]$
Recruitment Platform: CloudResearch $1.95$ $5.04^{*}$ $-6.02^{*}$ $-3.70^{*}$
$[-1.10;5.00]$ $[1.70;8.37]$ $[-9.65;-2.40]$ $[-7.05;-0.35]$
Num. obs. $2501$ $2427$ $2480$ $2458$
Num. groups. $463$ $465$ $458$ $473$
Null hypothesis value outside the confidence interval.
Table 14: Full population – Emotions: Increased Exposure Experiment, In-Feed. Regression tables of the analysis of the emotion outcomes for the in-feed survey responses in the increased exposure experiment.
Reduced Exposure, Post-Experiment
Angry Sad Excited Calm
(Intercept) $1.14^{*}$ $1.08^{*}$ $1.19^{*}$ $2.09^{*}$
$[0.97;1.32]$ $[0.91;1.24]$ $[1.01;1.37]$ $[1.88;2.31]$
Condition: Treatment $-0.08$ $-0.00$ $0.04$ $0.04$
$[-0.21;0.04]$ $[-0.12;0.12]$ $[-0.09;0.18]$ $[-0.10;0.17]$
Pre-survey Answer $0.45^{*}$ $0.44^{*}$ $0.54^{*}$ $0.41^{*}$
$[0.39;0.52]$ $[0.38;0.51]$ $[0.48;0.60]$ $[0.35;0.47]$
Recruitment Platform: CloudResearch $0.11$ $0.09$ $-0.08$ $-0.04$
$[-0.03;0.24]$ $[-0.04;0.21]$ $[-0.22;0.06]$ $[-0.18;0.11]$
Num. obs. $727$ $727$ $727$ $727$
Null hypothesis value outside the confidence interval.
Table 15: Full population – Emotions: Reduced Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses in the reduced exposure experiment.
Increased Exposure, Post-Experiment
Angry Sad Excited Calm
(Intercept) $0.90^{*}$ $0.87^{*}$ $1.19^{*}$ $2.05^{*}$
$[0.70;1.11]$ $[0.66;1.09]$ $[0.97;1.41]$ $[1.78;2.31]$
Condition: Treatment $-0.02$ $0.08$ $0.09$ $-0.07$
$[-0.17;0.13]$ $[-0.07;0.23]$ $[-0.06;0.25]$ $[-0.23;0.09]$
Pre-survey Answer $0.58^{*}$ $0.55^{*}$ $0.56^{*}$ $0.43^{*}$
$[0.50;0.65]$ $[0.47;0.63]$ $[0.49;0.63]$ $[0.36;0.50]$
Recruitment Platform: CloudResearch $0.13$ $0.22^{*}$ $-0.27^{*}$ $-0.12$
$[-0.03;0.28]$ $[0.06;0.37]$ $[-0.43;-0.11]$ $[-0.28;0.05]$
Num. obs. $528$ $528$ $528$ $528$
Null hypothesis value outside the confidence interval.
Table 16: Full population – Emotions: Increased Exposure Experiment, Post-Experiment. Regression tables of the analysis of the emotion outcomes for the post-experiment survey responses in the increased exposure experiment.
E.3 Political attitude shifts
As preregistered, we report the effect of the intervention on the seven remaining attitudes we used to identify the AAPA posts. We find that Reduced Exposure to AAPA content reduces the Biased Evaluation of Politicized Facts while increasing the exposure increases the Opposition to Bipartisan Cooperation (Fig. 16). However, none of these effects are significant after adjusting for multiple hypotheses testing as preregistered (Sec. D).
Figure 16: Treatment effects on anti-democratic attitudes estimated by regressing the participants’ post-survey responses on the treatment indicator, their pre-survey response to the same question, and an indicator of the platform they were recruited on. The p values are adjusted for multiple hypotheses testing as preregistered.
Figure 17: Each value represents the estimated effect from a separate regression model. The dependent variable is the political feeling response in the post-experiment survey, controlling for both the pre-experiment survey response and the fraction of political posts viewed over a ten-day period for at least one second, which expressed support for a particular factor (i.e., $post\_experiment\_rating\sim pre\_experiment\_rating+fraction\_v_{n}$, where $v_{n}$ is a specific factor such as partisan animosity or social distrust). The coefficient represents the estimated change (on a 0-100 scale) in feelings toward people from the opposite party when 100% of the viewed political posts expressed support for the corresponding factor ($v_{n}$). All participants are included in the analysis (N=1,256).
E.4 Engagement
During the baseline period, we observed that political posts enjoyed a higher reshare rate: 0.677%, compared to 0.484% for non-political posts (a relative increase of 39.7%; U=714,022; P$=$0.046). Similarly, political content had a higher favorite rate, with an average engagement rate for posts viewed for at least one second of 5.07% vs. 4.29% for non-political posts (U=645,266; P$=$0.001). We observed that AAPA posts, while shared more frequently than non-AAPA content, did not exceed political non-AAPA posts in share rate (0.662% vs. 0.674%; U=668,945, P=0.003). Similarly, although AAPA content had a higher favorite rate than non-AAPA posts, it was favorited less than non-AAPA political posts (4.6% vs. 5.23%; U=711,396, P$<$0.001).
When examining the effects of our interventions, none of the differences reached statistical significance. In the Increased Exposure experiment, the average number of daily sessions per participant–defined as scrolling activities separated by at least one hour–was 1.65 in the control group and 1.53 in the treatment group (U=32,220; P=0.18). In the Reduced Exposure experiment, the control and treatment groups averaged 1.53 and 1.49 sessions per day, respectively (U=66,940; P=0.71). After controlling for the average daily sessions during the baseline period and the recruitment platform, the change in session count was not statistically significant in either experiment (Increased Exposure: -0.07 sessions, CI [-0.19, 0.04]; P=0.197; 95% Reduced Exposure: -0.04 sessions, CI [-0.13, 0.05] P=0.381; 95%). The differences in the time spent on the platform across conditions also do not reach statistical significance. In the Increased Exposure experiment, participants in the control condition spent on average of 44.4 minutes vs. 41.3 minutes in the treatment condition (U=37,545; P=0.12), while in the Reduced Exposure, they spent 46.8 minutes in the control and 42.2 minutes in the treatment condition (U=70,678; P=0.10).
Similarly, repost rates during the experimental period did not differ significantly between treatment and control groups. In the Increased Exposure experiment, the repost rate was 0.52% in the control group and 1.07% in the treatment group (U=30,859; P=0.83). In the Reduced Exposure experiment, the repost rates were 0.74% for the control group and 0.46% for the treatment group. When controlling for user baseline averages and the platform using regression analysis (Fig. 18), the changes in repost rates remained non-significant (Increased Exposure: increase of 0.27%, 95% CI [-0.002, 0.008]; P=0.29; Reduced Exposure: reduction of -0.06%, 95% CI [-0.002, 0.001] P=0.503;).
The favorite rates during the experimental period also showed no significant differences. The overall favorite rate during the baseline was 4.73%. Controlling for user baseline and platform (Fig. 19), the favorite rate changed by 0.14% in the Increased Exposure group (95% CI [-0.009, 0.012]; P=0.796) and decreased by 0.29% in the Reduced Exposure group (95% CI [-0.010, 0.004]; P=0.43).
Finally, we observed no significant differences between treatment and control groups in terms of the number of new posts or replies during the experimental period.
Figure 18: Treatment effect on repost rate. Coefficient of a linear regression predicting participants’ repost rate during the experiment based on their baseline repost rate and experimental condition (treatment or control). Error bars represent 95% confidence intervals.
Figure 19: Treatment effect on favorite rate. Coefficient of a linear regression predicting participants’ favorite rate during the experiment based on their baseline favorite rate and experimental condition (treatment or control). Error bars represent 95% confidence intervals.
E.5 Heterogeneous Treatment Effects
Fig. 20 and Fig. 21 show the heterogeneous treatment effects on affective polarization in the two experiments. We use the same regression specifications as in the main analyses but interact the treatment with the moderator. As in the covariate balance and attrition tests, we coarsened the education, income, and SES ladder covariates to avoid small categories. We consider only a subset of covariates and split them into primary and secondary moderators as preregistered. After adjusting the p-values for multiple hypotheses testing, none of the covariates reach statistical significance.
Figure 20: Heterogeneous treatment effect estimates of reducing and increasing exposure to AAPA content on affective polarization based on the in-feed measurements. The p values are adjusted for multiple hypotheses testing as preregistered.
Figure 21: Heterogeneous treatment effect estimates of reducing and increasing exposure to AAPA content on affective polarization based on the post-experiment measurements. The p values are adjusted for multiple hypotheses testing as preregistered.
E.6 Qualitative feedback on the impact of the experiment
At the end of the study, in the post-survey, we asked participants: “Did the Chrome extension affect your experience on Twitter/X in any way?” This question was open-ended, and 1,218 participants (96.9%) provided a non-empty response. Most answers were short and indicated no impact, such as “No” or “Not that I could tell.” We individually coded each response with the support of AI (GPT-4), providing the original survey question for context and asking whether the participant reported any impact on their experience on Twitter/X.
The majority of participants (overall: 74.2%; increase, control: 73.7%; increase, treatment: 72.9%; reduce, control: 76.4%; reduce, treatment: 73.4%) explicitly stated that they did not notice any impact on their experience. We manually reviewed a random sample of 100 responses and found complete agreement with GPT-4’s annotations. Among the remaining responses, we manually coded 10.3% as reporting a “positive experience,” often for reasons unrelated to the extension, such as: “Twitter seemed better organized,” “It made me feel informed,” and “It helped me explore the social network better.” Additionally, 4.1% reported observing more political content; however, this observation was uniformly distributed across conditions, likely reflecting an overall increase in political content during the election period. Finally, 2.8% of participants mentioned noticing a slowdown in the website’s performance. The remaining answers mention details unrelated to the assigned condition, such as “It made me use it on my laptop more, which was a bit annoying” or “Having to use the For You page I noticed a lot of tweets about AI.”
To investigate this further, we adopted a quantitative approach by embedding participant responses using a pre-trained BERT-based model (all-MiniLM-L6-v2). We then predicted the assigned experiment using logistic regression with L2 regularization, employing 5-fold cross-validation. The classification achieved an F1 score of 53.9% ($\pm$ 4.7%) on a balanced dataset. Similarly, we predicted the experimental condition (treatment or control) for each experiment. For the reduced experiment, the F1 score was 47.4% ($\pm$ 6.8%), and for the increased experiment, it was 50.5% ($\pm$ 4.7%).
These results suggest that the experimental condition did not significantly impact participants’ experience, as they did not perceive or report notable differences.